Week of 230417

WLCG Operations Call details

  • The connection details for remote participation are provided on this agenda page.

General Information

  • The purpose of the meeting is:
    • to report significant operational issues (i.e. issues which can or did degrade experiment or site operations) which are ongoing or were resolved after the previous meeting;
    • to announce or schedule interventions at Tier-1 sites;
    • to inform about recent or upcoming changes in the experiment activities or systems having a visible impact on sites;
    • to provide important news about the middleware;
    • to communicate any other information considered interesting for WLCG operations.
  • The meeting should run from 15:00 Geneva time until 15:20, exceptionally to 15:30.
  • The SCOD rota for the next few weeks is at ScodRota
  • Whenever a particular topic needs to be discussed at the operations meeting requiring information from sites or experiments, it is highly recommended to announce it by email to the wlcg-scod list (at cern.ch) to allow the SCOD to make sure that the relevant parties have the time to collect the required information, or invite the right people at the meeting.

Best practices for scheduled downtimes

Monday

Attendance:

  • remote: Andrew P (NL-T1), Andrew (TRIUMF), Borja (Chair, Monitoring), Daniele (CNAF), Darren (RAL), David (IN2P3), David (FNAL), Doug (BNL), Ivan (ATLAS), Julia (WLCG), Priscilla (ATLAS), Xavier (KIT)

Experiments round table:

  • CMS reports (raw view) -
    • Note: Today there will likely nobody be available from CMS (due to collaboration meeting and other commitments)
    • Started to work with sites (running XRootD storage) to configure Tokens
      • Communication via GGUS tickets

  • ALICE
    • NTR

Sites / Services round table:

  • ASGC: NC
  • BNL: NTR
  • CNAF: NTR
  • EGI: NC
  • FNAL: dCache tape instance down tomorrow 8AM-5PM FNAL time for upgrade
  • IN2P3: 2 disks (178 TB) lost on last Saturday on our XRootD service for ALICE. Files are unavailable until we receive the new disks.
  • JINR: NTR
  • KISTI: NC
  • KIT: NTR
  • NDGF: NC
  • NL-T1: Nikhef compute and storage will be down between 17:00 and 22:00 CET today due to a network upgrade. See GOCDB:33840
  • NRC-KI: NC
  • OSG: NC
  • PIC: NC
  • RAL: Echo is going into a scheduled down-time (20/4/23) to allow patching and rebooting. GOCDB:33844
  • TRIUMF: NTR

  • CERN computing services: NC
  • CERN storage services: NC
  • CERN databases: NC
  • GGUS: NTR
  • Monitoring: NTR
  • Middleware: NTR
  • Networks: NC
  • Security: NC

AOB:

Edit | Attach | Watch | Print version | History: r12 < r11 < r10 < r9 < r8 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r12 - 2023-04-17 - NikolayVoytishin
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LCG All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2023 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
or Ideas, requests, problems regarding TWiki? use Discourse or Send feedback